# dire_short.bib

### v2

@@ Abstract
Summary: DiRe – JAX is a new dimensionality reduction toolkit designed to address some of the challenges faced by traditional methods like UMAP and tSNE such as loss of global structure and computational efficiency. Built on the JAX framework, DiRe leverages modern hardware acceleration to provide an efficient, scalable, and interpretable solution for visualizing complex data structures, and for quantitative analysis of lower–dimensional embeddings. The toolkit shows considerable promise in preserving both local and global structures within the data as compared to state–of–the–art UMAP and tSNE implementations. This makes it suitable for a wide range of applications in machine learning, bio–informatics, and data science.
1 Statement of Need
@@
However, these methods often struggle with scalability, interpretability, and preservation of global data structures. UMAP, while fast and scalable, may overemphasize local structures at the expense of global data relationships [pachter]. And tSNE, while known for producing high–quality visualizations, may be computationally expensive and sensitive to hyperparameter tuning [kobak2019art].
DiRe–JAX addresses these challenges by offering a scalable solution that balances the preservation of both local and global structures. Leveraging the JAX framework allows DiRe–JAX to efficiently handle large datasets by utilizing GPU acceleration, making it significantly faster than non–GPU implementations without compromising on the quality of the embeddings.
@@ 2 Main methods
    Capturing dataset topology: Create the kNN graph of $X$, say $\Gamma$, for a given number of neighbors $k$ ($=$ n_neighbors) by calling make_knn_adjacency. This step uses FAISS [faiss] which is the most efficient kNN library to the best of authors’ knowledge.
@@ 2.1 Random Projection embedding
This embedding is based on the following famous lemma (in fact, a rather useful theorem) of Johnson and Lindenstrauss [JL].
@@
Given $10<\epsilon<$ 1 and an integer $n$, let $X$ be a set of $n$ points in $\mathbb{R}^{d}$. For a random linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ where $k=O\left(\frac{\log n}{\epsilon^{2}}\right)$, with high probability, for all $u,v\in X$,
@@ 2.2 Principal Component Analysis embedding
If we believe that most of the topological features of $X$ are expressed in dimensions up to $k$, while the remaining dimensions are negligible, we then contend that the persistence diagrams $D(X)$ of $X$ in the first $k$ dimensions may be replaced by those of $\widehat{X}$. The actual $k$–dimensional image of $\widehat{X}$ is $\widehat{X}\,W_{k}$. Thus, in the first $k$ dimensions we put $D(X):=D(\widehat{X}\,W_{k})$. We naturally have this equality if $X$ belongs to an affine $k$–dimensional subspace of $\mathbb{R}^{d}$.
Then, according to [chazal-persistence], for the persistence diagrams we have that the bottleneck distance between the persistence diagrams of $X$ and $X_{k}$ or, equivalently, $X_{k}$ satisfies
@@ 2.3 Spectral Laplacian embedding
Yet another popular embedding method which may or may not preserve the internal structure of the data. It is often geared towards manifold learning of the local data structure in order to deal with non–linearity, even though unwrapping non–linear data in lower dimensions may or may not be beneficial for its subsequent analysis. The same applies to diffusion maps. Since the $kNN$–graph $\Gamma$ of $X$ is already available, the Laplacian embedding can be computed with standard methods.
@@ 2.4 Force–directed layout
After the initial embedding, DiRe–JAX applies an iterative force–directed layout in order to correlate the local structure of the embedding $Y$ with the local structure of the high–dimensional dataset $X$. The forces of attraction and repulsion are given by simple distributions modeled after the $t$–distributions of UMAP and tSNE.
@@
is fit to min_dist $=\delta$ and spread $=\sigma$ parameters, so that
@@
The attraction forces are defined between the point in $Y$ that are projections of $kNN$–neighbors in $\Gamma$, of magnitude
@@ 3.1.1 Persistence diagrams
Persistence diagrams are classical representations of the “birth–death” pairs in persistence homology. The reader may find more information in [ZC, zomorodian2005topology], while we give a brief description below.
@@
where we use $H_{k}(X)$ to denote the $k$–dimensional persistence homology group defined above.
@@
The usual metrics between two diagrams $D=D_{k}(X)$ and $D^{\prime}=D_{k}(Y)$ are the bottleneck and Wasserstein distances, cf. [persim-docs].
@@ 3.1.2 Betti curves
Betti curves mark the progression of the persistence Betti numbers with the filtration level increase. This representation was first studied in detail in [Giusti-et-al]. Let $X$ be a point cloud and $\mathrm{Rips}_{t}(X)$ be its Rips complex of level $t\geq 0$. Then the $k$–th Betti curve is
@@
The most common are the Dynamic Time Warp (DTW) distance [salvador-fastdtw, fastdtw-docs], Time Warp Edit Distance (TWED) [marteau-twed, pytwed-docs], and Earth Mover’s Distance (EMD) [flamary-pot, ot-docs].
@@ 3.1.3 Global structure preservation
Let $X\in\mathbb{R}^{n\times d}$ be a set of $n$ vectors in $\mathbb{R}^{d}$ with $kNN$–graph $\Gamma$. Let $Y=f(X)\in\mathbb{R}^{n\times k}$ be its embedding. We shall measure the most common distances between the corresponding persistence diagrams $D^{H}_{k}=D_{k}(X)$ and $D^{L}_{k}=D_{k}(Y)$ of higher–dimensional data and its lower–dimensional embedding, respectively. Namely, we will use the bottleneck $d_{b}(D^{H}_{k},D^{L}_{k})$ and Wasserstein distances $d_{W}(D^{H}_{k},D^{L}_{k})$ for $k=0,1$, as our embedding will be $2$–dimensional in all practical examples.
@@
Again, we compute their normalized versions: averaged over the number of subsampled points. Also, the curves are first normalized to the unit mass for the EMD computation, and later on the distance is rescaled back. Indeed, we are first interested in how much the mass has to be moved on the relative scale, and the rescaling does not change the essential topology of the mass distribution. If $\tau$ is our normalizing factor, we thus compute $\max\{\tau,\tau^{-1}\}\cdot d_{EMD}(\beta^{H},\tau\cdot\beta^{L})$.
@@ 3.2.1 Embedding stress
Let $X\in\mathbb{R}^{n\times d}$ be a set of $n$ vectors in $\mathbb{R}^{d}$ with $kNN$–graph $\Gamma$. Let $Y=f(X)\in\mathbb{R}^{n\times k}$ be its embedding. For each edge $e=(x,y)$ in $\Gamma$ we have its local embedding stress measured by
@@ 4 General workflow
5 Benchmarks
A collection of benchmarks is available in our GitHub repository [dire-github], some of which are given below. We shall concentrate on the main differences between DiRe–JAX, tSNE, and UMAP. Both tSNE and UMAP were implemented by Rapids.ai, and these are the GPU–based implementations that we use for benchmarking. We also use the original UMAP implementation by McInnes et al.
@@ 5 Benchmarks
5.1 Dataset: Blobs
The 2D embeddings of the blobs dataset with 12 blobs having $10$K points $1000$–dimensional points in total are given in Fig. 1. We can see that despite the correct parameter cuML tSNE produces strangely erroneous output. We can also observe from the pictures that cuML UMAP does not produce very smooth round blobs, while the original UMAP does. Also, DiRE–JAX produces a visually pleasing embedding with round blobs having smooth boundaries.
The average neighborhood preservation score is very low for all four methods (less than $1$%). This can be easily explained with the following picture: a point cloud in 2D with the points $X_{k}=\{x_{n}=(1/n,n)\,:\,n=1,\ldots,k\}$ has $x_{k}$ as the largest distance from the origin, yet $x_{k}$ projects to the closest point to 00 on the horizontal axis.
(a) DiRe – JAX
(b) cuML tSNE
(c) cuML UMAP
(d) Original UMAP
Figure 2: ”MNIST Digits” dataset embeddings
@@ 5.2 Dataset: MNIST Digits
The key metrics are given in Appendix 8, which shows that tSNE performs much better on persistence homology global measures. The difference is drastic in dimension 00, so that it appears tSNE is winning in this case in plotting the actual topological layout of MNIST digits.
@@
5.3 Dataset: Disk Uniform
Here we embed a 2D disk with points uniformly distributed in it. In the earnest, one would expect that uniformity is preserved, especially by UMAP, as in Uniform Manifold Approximation and Projection, and as it purports to make the distribution more uniform by rescaling [umap]. However, except for DiRe–JAX none of the methods visibly produces anything close to uniformity. Even more prominently, cuML UMAP fails to produce a connected embedding for one of the major clusters.
A very similar picture arises when we try to embed a sphere and an ellipsoid in 3D with uniformly distributed points. The ellipsoid case is most egregious, as dilation in the principal axes leads to even more distortion of uniformity in the embeddings (both 2D and 3D). See the GitHub repository for more datasets [dire-github].
@@ 5.5 Dataset: Levine 13
More precisely, cuML tSNE outperforms all other methods on most persistence metrics for dimension 00, except the bottleneck distance.
@@ 5.6 Dataset: Levine 32
6 Future work
The following points are important for future work and the subsequent releases of DiRe–JAX.
  1. 1.
    Adding more quantitative measures for analysis and optimization, as well as hyperparameter optimization functionality (suitable for smaller point clouds).
  2. 2.
    Scaling the algorithm past GPU memory limitations. This cannot be done simply by batch processing, as not all mutual positions of the data points will be accessible within one batch in order to create the correct $kNN$–graph.
  3. 3.
    Producing a pure TPU version: for now this can be done by running FAISS on CPU, and installing the TPU version of JAX, but that is obviously not as efficient as running kNN/ANN on the TPU directly. The freely available ANN TPU–based package annax [annax] is not even nearly as efficient as FAISS, and in our experiments were not able to handle a dataset with $10$K points of dimension $1000$ by exhausting all GPU memory.
Acknowledgments
[…]
7 Blobs: Dataset Metrics
(a) Runtime (seconds)
(b) Embedding stress
(c) Neighborhood preservation
(d) Context preservation
(e) Persistence homology preservation: dimension 00 (relative to DiRe–JAX baseline)
(f) Persistence homology preservation: dimension $1$ (relative to DiRe–JAX baseline)
Figure 7: ”Blobs” dataset metrics comparison
8 MNIST Digits: Dataset Metrics
(a) Runtime (seconds)
(b) Embedding stress
(c) Neighborhood preservation
(d) Context preservation
(e) Persistence homology preservation: dimension 00 (relative to DiRe–JAX baseline)
(f) Persistence homology preservation: dimension $1$ (relative to DiRe–JAX baseline)
Figure 8: ”MNIST Digits” dataset metrics comparison
9 Disk Uniform: Dataset Metrics
(a) Runtime (seconds)
(b) Embedding stress
(c) Neighborhood preservation
(d) Persistence homology preservation: dimension 00 (relative to DiRe–JAX baseline)
(e) Persistence homology preservation: dimension $1$ (relative to DiRe–JAX baseline)
Figure 9: ”Disk Uniform” dataset metrics comparison
10 Half–moons: Dataset Metrics
(a) Runtime (seconds)
(b) Embedding stress
(c) Neighborhood preservation
(d) Context preservation
(e) Persistence homology preservation: dimension 00 (relative to DiRe–JAX baseline)
(f) Persistence homology preservation: dimension $1$ (relative to DiRe–JAX baseline)
Figure 10: ”Half–moons” dataset metrics comparison
11 Levine 13: Dataset Metrics
(a) Runtime (seconds)
(b) Embedding stress
(c) Neighborhood preservation
(d) Context preservation
(e) Persistence homology preservation: dimension 00 (relative to DiRe–JAX baseline)
(f) Persistence homology preservation: dimension $1$ (relative to DiRe–JAX baseline)
Figure 11: ”Levine 13” dataset metrics comparison
12 Levine 32: Dataset Metrics
(a) Runtime (seconds)
(b) Embedding stress
(c) Neighborhood preservation
(d) Context preservation
(e) Persistence homology preservation: dimension 00 (relative to DiRe–JAX baseline)
(f) Persistence homology preservation: dimension $1$ (relative to DiRe–JAX baseline)
Figure 12: ”Levine 32” dataset metrics comparison

### v3

# Dimensionality reduction for homological stability and global structure preservation

@@ Abstract
We propose a new dimensionality reduction toolkit designed to address some of the challenges faced by traditional methods like UMAP and tSNE such as loss of global structure and computational efficiency. Built on the JAX framework, DiRe leverages modern hardware acceleration to provide an efficient, scalable, and interpretable solution for visualizing complex data structures, and for quantitative analysis of lower–dimensional embeddings. The toolkit shows considerable promise in preserving both local and global structures within the data as compared to state–of–the–art UMAP and tSNE implementations. This makes it suitable for a wide range of applications in machine learning, bio–informatics, and data science.
1 Introduction
@@
However, these methods often struggle with scalability, interpretability, and preservation of global data structures. UMAP, while fast and scalable, may overemphasize local structures at the expense of global data relationships [CP23]. And tSNE, while known for producing high–quality visualizations, may be computationally expensive and sensitive to hyperparameter tuning [KB19].
DiRe–JAX addresses these challenges by offering a scalable solution that balances the preservation of both local and global structures. Leveraging the JAX framework allows DiRe–JAX to efficiently handle large datasets by utilizing GPU / TPU acceleration, making it significantly faster than CPU–based implementations without compromising on the quality of the embeddings.
While the software was reviewed and published in [KR25], the required theoretical background and detailed benchmark analysis constitute the main subjects of this paper.
@@ 2 Main methods
    Capturing dataset topology: Create the kNN graph of $X$, say $\Gamma$, for a given number of neighbors $k$ (== n_neighbors) by calling make_knn_adjacency. This step uses a JAX kernel specifically developed by the authors in order to perform the computation int the CPU, GPU or TPU settings. Other libraries like FAISS [Dou+24] may also be used, although they do not have the same level of hardware universality.
@@ 2.1 Random Projection embedding
This embedding is based on the following famous lemma (in fact, a rather useful theorem) of Johnson and Lindenstrauss [JL84].
@@
Given $0<\epsilon<1$ and an integer $n$, let $X$ be a set of $n$ points in $\mathbb{R}^{d}$. For a random linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ where $k=O\left(\frac{\log n}{\epsilon^{2}}\right)$, with high probability, for all $u,v\in X$,
@@ 2.2 Principal Component Analysis embedding
If we believe that most of the topological features of $X$ are expressed in dimensions up to $k$, while the remaining dimensions are negligible, we then contend that the persistence diagrams $D(X)$ of $X$ in the first $k$ dimensions may be replaced by those of $\widehat{X}$. The actual kk–dimensional image of $\widehat{X}$ is $\widehat{X}\,W_{k}$. Thus, in the first $k$ dimensions we put $D(X):=D(\widehat{X}\,W_{k})$. We naturally have this equality if $X$ belongs to an affine kk–dimensional subspace of $\mathbb{R}^{d}$.
Then, according to [CSO14], for the persistence diagrams we have that the bottleneck distance between the persistence diagrams of $X$ and $X_{k}$ or, equivalently, $X_{k}$ satisfies
@@ 2.3 Spectral Laplacian embedding
Yet another popular embedding method which may or may not preserve the internal structure of the data. It is often geared towards manifold learning of the local data structure in order to deal with non–linearity, even though unwrapping non–linear data in lower dimensions may or may not be beneficial for its subsequent analysis. The same applies to diffusion maps. Since the kNNkNN–graph $\Gamma$ of $X$ is already available, the Laplacian embedding can be computed with standard methods.
@@ 2.4 Force–directed layout
After the initial embedding, DiRe–JAX applies an iterative force–directed layout in order to correlate the local structure of the embedding $Y$ with the local structure of the high–dimensional dataset $X$. The forces of attraction and repulsion are given by simple distributions modeled after the tt–distributions of UMAP and tSNE.
@@
is fit to min_dist $\delta$ and spread $\sigma$ parameters, so that
@@
The attraction forces are defined between the point in $Y$ that are projections of kNNkNN–neighbors in $\Gamma$, of magnitude
@@ 3.1.1 Persistence diagrams
Persistence diagrams are classical representations of the “birth–death” pairs in persistence homology. The reader may find more information in [ZC05, Zom05], while we give a brief description below.
@@
where we use $H_{k}(X)$ to denote the kk–dimensional persistence homology group defined above.
@@
The usual metrics between two diagrams $D=D_{k}(X)$ and $D^{\prime}=D_{k}(Y)$ are the bottleneck and Wasserstein distances, cf. [Per24].
@@ 3.1.2 Betti curves
Betti curves mark the progression of the persistence Betti numbers with the filtration level increase. This representation was first studied in detail in [GGB16]. Let $X$ be a point cloud and $\mathrm{Rips}_{t}(X)$ be its Rips complex of level $t\geq 0$. Then the kk–th Betti curve is
@@
The most common are the Dynamic Time Warp (DTW) distance [SC07, Dev23], Time Warp Edit Distance (TWED) [Mar09, Zum21], and Earth Mover’s Distance (EMD) [Fla+21, POT25].
@@ 3.1.3 Global structure preservation
Let $X\in\mathbb{R}^{n\times d}$ be a set of $n$ vectors in $\mathbb{R}^{d}$ with kNNkNN–graph $\Gamma$. Let $Y=f(X)\in\mathbb{R}^{n\times k}$ be its embedding. We shall measure the most common distances between the corresponding persistence diagrams $D^{H}_{k}=D_{k}(X)$ and $D^{L}_{k}=D_{k}(Y)$ of higher–dimensional data and its lower–dimensional embedding, respectively. Namely, we will use the bottleneck $d_{b}(D^{H}_{k},D^{L}_{k})$ and Wasserstein distances $d_{W}(D^{H}_{k},D^{L}_{k})$ for $k=0,1$, as our embedding will be 22–dimensional in all practical examples.
@@
Again, we compute their normalized versions: averaged over the number of subsampled points. Also, the curves are first normalized to the unit mass for the EMD computation, and later on the distance is rescaled back. Indeed, we are first interested in how much mass has to be moved on the relative scale, and the rescaling does not change the essential topology of the mass distribution. If $\tau$ is our normalizing factor, we thus compute $\max\{\tau,\tau^{-1}\}\cdot d_{EMD}(\beta^{H},\tau\cdot\beta^{L})$.
@@ 3.2.1 Embedding stress
Let $X\in\mathbb{R}^{n\times d}$ be a set of $n$ vectors in $\mathbb{R}^{d}$ with kNNkNN–graph $\Gamma$. Let $Y=f(X)\in\mathbb{R}^{n\times k}$ be its embedding. For each edge $e=(x,y)$ in $\Gamma$ we have its local embedding stress measured by
@@ 4 General workflow
5 Benchmarking
A collection of benchmarks is available in our GitHub repository [KR25a], some of which are given below. We shall concentrate on the main differences between DiRe–JAX, tSNE, and UMAP. Both tSNE and UMAP were implemented by Rapids.ai [Dev25], and these are the GPU–based implementations that we use for benchmarking. We also use the original UMAP implementation by McInnes et al. [Dev25a].
5.1 Dataset: Blobs
The 2D embeddings of the blobs dataset with 12 blobs having 1010K points 10001000–dimensional points in total are given in Fig. 1. We can see that despite the correct parameter cuML tSNE produces strangely erroneous output. We can also observe from the pictures that cuML UMAP does not produce very smooth round blobs, while the original UMAP does. Also, DiRE–JAX produces a visually pleasing embedding with round blobs having smooth boundaries.
The average neighborhood preservation score is very low for all four methods (less than 11%). This can be easily explained with the following picture: a point cloud in 2D with the points $X_{k}=\{x_{n}=(1/n,n)\,:\,n=1,\ldots,k\}$ has $x_{k}$ as the largest distance from the origin, yet $x_{k}$ projects to the closest point to 0 on the horizontal axis.
@@ 5.2 Dataset: MNIST Digits
The key metrics are given in Appendix 8, which shows that tSNE performs much better on persistence homology global measures. The difference is drastic in dimension 0, so that it appears tSNE is winning in this case in plotting the actual topological layout of MNIST digits.
@@
Figure 2: ”MNIST Digits” dataset embeddings
5.3 Dataset: Disk Uniform
Here we embed a 2D disk with points uniformly distributed in it. In the earnest, one would expect that uniformity is preserved, especially by UMAP, as in Uniform Manifold Approximation and Projection, and as it purports to make the distribution more uniform by rescaling [MHM18]. However, except for DiRe–JAX none of the methods visibly produces anything close to uniformity. Even more prominently, cuML UMAP fails to produce a connected embedding for one of the major clusters.
A very similar picture arises when we try to embed a sphere and an ellipsoid in 3D with uniformly distributed points. The ellipsoid case is most egregious, as dilation in the principal axes leads to even more distortion of uniformity in the embeddings (both 2D and 3D). See the GitHub repository for more datasets [KR25a].
(a) DiRe – JAX
(b) cuML tSNE
(c) cuML UMAP
(d) Original UMAP
@@ 5.5 Dataset: Levine 13
More precisely, cuML tSNE outperforms all other methods on most persistence metrics for dimension 0, except the bottleneck distance.
@@ 5.6 Dataset: Levine 32
6 Conclusions
A new dimensionality reduction algorithm, called DiRe, has been developed and its main properties established theoretically: the most important of them is reasonable preservation of the dataset’s homological structure. The latter is quantified by persistence homology and expressed via persistence diagrams or Betti curves. The benchmarks and metrics provided lead to a conclusive practical demonstration of the claim that in most cases DiRe preserves the global dataset structure better than tSNE or UMAP.
The software used in the numerical experiments is available from GitHub [KR25a], and has been peer reviewed with a published version of record [KR25]. The current stable version of DiRe is also available as a Python package distribution, and the most recent development version is accessible on GitHub together with detailed API documentation.
The reader can reproduce the computational results of this paper by accessing GitHub [KR25a] and running the benchmarking Jupyter notebook in Google Colab. In order to avoid lengthy runtime, GPU / TPU acceleration is required.